OpenStudy (anonymous):

What is the slant height of a square pyramid that has a surface area of 525 square meters and a side length of 15 meters?

OpenStudy (anonymous):

\[\large \color{red}{S.A. = (Side)^2 + 2 \times l \times s}\] S.A. is Surface Area of Square Pyramid: L is slant height and s is Side

OpenStudy (anonymous):

So: \[\large 525 = (15)^2 + 2 \times l \times (15)\] \[\large 525 = 225 + 30l\] Subtract both the sides by 225 first and then divide both side by 30 to get Slant Height(l)..